solve

1.d2y/dx2-2dy/dx+5y=extan2x

2.dy/dx+3y=3x2e-3x

1
Expert's answer
2021-07-06T14:36:58-0400

Answer:-

1."{d^2y\over dx^2}-2{dy\over dx}+5y=e^xtan2x"

The differential equation has the form

"{d^2y\over dx^2}+p{dy\over dx}+qy=s" where "p=-2" , "q=5" and "s=-e^xtan2x"

Solve corresponding homogenous linear equation

"{d^2y\over dx^2}+p{dy\over dx}+qy=0"

Its characteristic equation is given as

"q+(k^2+kp)=0\implies5+(k^2-2k)=0 \implies k^2-2k+5=0"

"\therefore" "k_1=1-2i" and "k_2=1+2i"

"\implies" "y=C_1e^{k_1x}+C_2e^{k_2x}=C_1e^{(1-2i)x}+C_2e^{(1+2i)x}"

Solve the inhomogeneous equation

"{d^2y\over dx^2}+p{dy\over dx}+qy=s"

By use of variation of parameters method, suppose that "C_1" and "C_2" are functions of "x"

"\therefore" The general solution is

"y=C_1(x)e^{(1-2i)x}+C_2(x)e^{(1+2i)x}"

To find "C_1(x)" and "C_2 (x)" ,by the method of variation of parameters ,we find the solution from the system:

"y_1(x){d\over dx}C_1(x)+y_2(x){d\over dx}C_2(x)=0"


"{d\over dx}C_1(x){d\over dx}y_1(x)+{d\over dx}C_2(x){d\over dx}y_2(x)=f(x)"


Where "y_1(x)" and "y_2(x)" are linearly independent particular solutions of linear ordinary differential Equations

"y_1(x)=e^{x(1-2i)}" , "C_1=1" and "C_2=0"


"y_2(x)=e^{x(1+2i)}" , "C_1=0" and "C_2=1"

And the free term is given as

"f(x)=e^xtan(2x)"

So the system has the form:

"e^{x(1-2i)}{d\over dx}C_1(x)+e^{x(1+2i)}{d\over dx}C_2(x)=0"


"{d\over dx}C_1(x){d\over dx}e^{x(1-2i)}+{d\over dx}C_2(x){d\over dx}e^{x(1+2i)}=e^xtan(2x)"


solving the system gives

"{d\over dx}C_1(x)={ie^{2ixtan(2x)}\over 4} \implies C_1(x)=\int{ie^{2ixtan(2x)}\over 4}dx"


"{d\over dx}C_2(x)={ie^{-2ixtan(2x)}\over 4}\implies C_2(x)=-\int {ie^{-2ixtan(2x)}\over 4}dx"

Evaluate "\int{ie^{2ixtan(2x)}\over 4}dx={log(e^{2ix}-i)\over 8}-{log(e^{2ix}+i)\over 8}-{ie^{2ix}\over 8}"


"-\int{ie^{-2ixtan(2x)}\over 4}dx={log(e^{2ix}-i)\over 8}-{log(e^{2ix}+i)\over 8}+{ie^{2ix}\over 8}"

Substitute "C_1(x)" and "C_2(x)" to


"y=C_1(x)e^{(1-2i)x}+C_2(x)e^{(1+2i)x}"


"\implies y=({log(e^{2ix}-i)\over 8}-{log(e^{2ix}+i)\over 8}-{ie^{2ix}\over 8})e^{(1-2i)x}+({log(e^{2ix}-i)\over 8}-{log(e^{2ix}+i)\over 8}+{ie^{2ix}\over 8})e^{(1+2i)x}"


"\therefore y=(C_2sin(2x)+(C_1+{log(sin(2x)-1)\over 8}-{log(sin(2x)+1)\over 8})cos(2x)e^x"


2."{dy\over dx}+3y=3x^2e^{-3x}"

This differential equation has the form

"{dy\over dx}+p(x)y=Q(x)" ​​

Solve the correspondent linear homogenous equation

"{dy\over dx}+p(x)y=0" where "p(x)=3" and "Q(x)=3x^2e^{-3x}"

"\implies {dy\over dx}=-p(x)y \implies {dy\over y}=-p(x)dx \implies \int{1\over y}dy =-\int p(x)dx"

"\therefore log(|y|)=-\int p(x)dx \implies |y|=e^{-\int p(x)dx}"

"\implies y_1=e^{-\int p(x)dx}=e^{-\int 3dx}=e^{C_1-3x}"

and "y_2=-e^{-\int p(x)dx}=-e^{-\int 3dx}=-e^{C_2-3x}"

This leads to correspondent solution for any constant C not equal to zero

"y=Ce^{-3x}"

Now we should solve the inhomogeneous equation

"{dy\over dx}+p(x)y=Q(x)"

Use variation of parameters method by considering C as a function of x

"y=C(x)e^{-3x}"

Apply it in the original equation using rules for product differentiation of composite functions derivative .We find that

"{d\over dx}C(x)=Q(x)e^{\int p(x)dx}"

Let use Q(x) and p(x) for this equation, we get the first order differential equation for C(x)

"{d\over dx}C(x)=3x^3 \implies C(x)=\int3x^2dx=x^3+c"

use C(x) at

"y=C(x)e^{-3x}"

"\therefore" "y=(C_1+x^3)e^{-3x}"